OpenStudy (anonymous):

If g(x)=4x+1 What is the value of g(-2)?

OpenStudy (solomonzelman):

g(x)=4x+1 if I wanted to find g(-3), I would plug in -3 for x into the function. g(-3)=4(-3)+1 g(-3)=-12+1 g(-3)=-11 can you find g(-2) ?

OpenStudy (anonymous):

g(-2)=4(-2)+1 g(-2)= -8+1 g(-2)= -7 Right?

OpenStudy (solomonzelman):

Yes, you are correct!

OpenStudy (solomonzelman):

Very good.

OpenStudy (anonymous):

Thanks! You deserve a medal :)

OpenStudy (solomonzelman):

Anytime... glad to help!